# OMNI-CLST: ERROR-AWARE CURRICULUM LEARNING WITH GUIDED SELECTIVE CHAIN-OF-THOUGHT FOR AUDIO QUESTION ANSWERING

### v2

@@ 2.3 Guided Thought Dropout Mechanism
During the GRPO stage, given an audio $a\in\mathcal{A}$ and a text query $q\in\mathcal{Q}$, we sample $N$ candidate responses $\{o_{1},o_{2},\ldots,o_{N}\}$ from the current policy $\pi_{old}$. Each response $o_{i}$ is evaluated with a reward function $r(\cdot)$, and the rewards are normalized by mean and standard deviation to define the advantage:
@@
We employ two reward functions during GRPO training. The first is the accuracy reward $R_{\text{acc}}$, which assigns $1$ if the predicted answer $o_{i}$ matches the ground truth $s_{i}$, and 0 otherwise:
@@ 3.2 Experimental Setup
During the SFT stage, we fine-tune the model with LoRA [20], using a learning rate of $1\mathrm{e}{-4}$, LoRA rank of $8$, LoRA $\alpha$ of $32$, for $2$ epochs with a batch size of $6$. In the GRPO stage, we sample training data according to model performance, with a ratio of pretrained model correct : SFT correct : SFT incorrect = 3:20:7, using 15k samples for one epoch. We train with a batch size of $1$, a temperature of $1.2$, 44 candidate responses per GRPO step, a learning rate of $1\mathrm{e}{-6}$, and $\beta=0.04$ in Equation 2. $\lambda_{\text{acc}}=2$ and $\lambda_{\text{f}}=1$ in Equation 5. All experiments are conducted on $4$ NVIDIA A100 GPUs (80GB).

### v3

@@ 2.3 Guided Thought Dropout Mechanism
During the GRPO stage, given an audio $a\in\mathcal{A}$ and a text query $q\in\mathcal{Q}$, we sample NN candidate responses $\{o_{1},o_{2},\ldots,o_{N}\}$ from the current policy $\pi_{old}$. Each response $o_{i}$ is evaluated with a reward function $r(\cdot)$, and the rewards are normalized by mean and standard deviation to define the advantage:
@@
We employ two reward functions during GRPO training. The first is the accuracy reward $R_{\text{acc}}$, which assigns 11 if the predicted answer $o_{i}$ matches the ground truth $s_{i}$, and 0 otherwise:
@@ 3.2 Experimental Setup
During the SFT stage, we fine-tune the model with LoRA [20], using a learning rate of $1\mathrm{e}{-4}$, LoRA rank of 88, LoRA $\alpha$ of 3232, for 22 epochs with a batch size of 66. In the GRPO stage, we sample training data according to model performance, with a ratio of pretrained model correct : SFT correct : SFT incorrect = 3:20:7, using 15k samples for one epoch. We train with a batch size of 11, a temperature of $1.2$, 44 candidate responses per GRPO step, a learning rate of $1\mathrm{e}{-6}$, and $\beta=0.04$ in Equation 2. $\lambda_{\text{acc}}=2$ and $\lambda_{\text{f}}=1$ in Equation 5. All experiments are conducted on 44 NVIDIA A100 GPUs (80GB).